# How to Learn a Star: Binary Classification with Starshaped Polyhedral Sets

### v1

@@ Abstract
We consider binary classification restricted to a class of continuous piecewise linear functions whose decision boundaries are (possibly nonconvex) starshaped polyhedral sets, supported on a fixed polyhedral simplicial fan. We investigate the expressivity of these function classes and describe the combinatorial and geometric structure of the loss landscape, most prominently the sublevel sets, for two loss-functions: the 0/1-loss (discrete loss) and an exponential loss function. In particular, we give explicit bounds on the VC dimension of this model, and concretely describe the sublevel sets of the discrete loss as chambers in a hyperplane arrangement. For the exponential loss, we give sufficient conditions for the optimum to be unique, and describe the geometry of the optimum when varying the rate parameter of the underlying exponential probability distribution.
@@ 1 Introduction
We study the problem of binary classification from a geometric and combinatorial perspective. Given a finite labeled data-set and a prescribed loss-function, we focus on characterizing the structure of those parameters that yield perfect classification – namely, the set of global minimizers of the loss function. More generally, we investigate the geometry and combinatorics of the entire loss landscape in parameter space. Understanding the geometry is central to analyze the behavior of learning algorithms, as, for example, the arrangement of critical points and the connectivity of minimizers influence optimization efficiency and generalization. Combinatorial structures, such as polyhedral decompositions, provide insights into how parameter spaces partition into regions of similar behavior. Naturally, these subdivisions interact with the subdivision into sets of classifiers which induce the same classification on the data, and is therefore intimately related to the VC-dimension (the Vapnik-Chervonenkis dimension) of binary classifiers (Vapnik and Chervonenkis, 1971).
@@
In order to make rigorous statements, we fix the function class used for classification to be a class which is suitable for the specific learning task. Fixing a large set of classifiers can lead to practical difficulties due to the complexity of the space of allowed classifiers, while a small set of classifiers may not be able to capture the nature of the underlying problem. A natural function class consists of those functions whose decision boundary – the geometric object separating the two classes – is the boundary of a convex polyhedron. Such function classes have been previously considered, for example, in Astorino and Gaudioso (2002); Manwani and Sastry (2010) and Kantchelian et al. (2014), where the optimal separating convex polyhedron is found through iteratively solving LPs, minimizing a logistic loss function and finding a large margin convex separator, respectively.
@@
While polyhedral classifiers form a well-structured function class, they are also highly restrictive. In particular, the region enclosed by the decision boundary is necessarily convex, which may not always align with the structure of the underlying classification problem. Maintaining the piecewise linear nature, but allowing non-convex functions, we consider a class of piecewise linear functions whose decision boundaries are (possibly nonconvex) star-shaped polyhedral sets, supported on a fixed polyhedral fan. Fixing the polyhedral fan makes this a tractable and learnable class of functions whose space of parameters exhibits nice geometric structures as we show. Considering these function classes generalizes the approach in Cevikalp and Triggs (2017), where kites are used for solving visual object detection and multi-class discrimination. With this choice of classifiers, we avoid certain undesirable combinatorial and geometric properties inherent to the parameter space of general piecewise linear functions (or equivalently, functions representable by ReLU neural networks) such as disconnectedness and the presence of local non-global minima of the $0/1$-loss in case of separable data (Brandenburg et al., 2024).
@@
We thus focus on classification with polyhedral starshaped sets with respect to two loss functions: the 0/1-loss (or discrete loss) allows us to study the parameter space in a combinatorial fashion using polyhedral methods. Additionally, we consider an exponential loss function, which is amenable to numerical optimization methods, and whose level sets carry convex-geometric structures.
@@ 1.1 Our contributions
We initiate the geometric study of the space of parameters defining the classification by polyhedral starshaped sets, and show that binary classifications with this model correspond to chambers in the data arrangement, a hyperplane arrangement within the parameter space. Investigating the expressivity of such starshaped polyhedral classifiers, we show that the VC dimension equals the number of rays in the fan.
@@
We describe the (sub)level sets of the aforementioned loss functions. We show that the sublevel sets of false positives and false negatives are starshaped sets in the parameter space, and, in case of perfect separability, sublevel sets with respect to the discrete loss are convex sets. For the exponential loss we show concavity, and consequently convexity holds for its superlevel sets, implying that an optimum can be found in polynomial time. We give sufficient conditions for the optimum to be unique, and describe the geometry of the optimum when varying the rate parameter of the underlying exponential probability distribution.
@@
While most of our results focus on starshaped sets with fixed origin, we also consider starshaped classifiers where we allow translation of the origin, and consider the translation vector as an extra parameter. For a fixed starshaped set and varying translation vectors, we show that the discrete loss is constant on chambers in an arrangement of stars, but the sublevel sets are generally no longer convex. While this setup combined with the exponential loss does in general not lead to a convex program, we show that the log-likelihood function is piecewise concave on the underlying data fan arrangement.
@@ 1.2 Limitations
Throughout the article, we assume a fixed simplicial fan as given. In practical applications, an appropriate fan suitable for the specific task needs to chosen prior to the analysis. We emphasize that this paper is a purely theoretical contribution. The presented framework has not been tested on large-scale synthetic or real-world data. Developing a systematic or heuristic approach for selecting parameters such as the simplicial fan, a translation vector or the rate parameter, is beyond the scope of this work.
@@ 2.1 Polyhedral geometry and stars
We begin by introducing essential notions from polyhedral geometry, and the class of classifiers we consider. For a thorough background on polyhedral geometry we refer the reader to Ziegler (2012).
@@
A set $S\subseteq\mathbb{R}^{d}$ is star-convex with respect to a center $\mathbf{o}\in S$ if for every $\mathbf{s}\in S$, the line segment $[\mathbf{o},\mathbf{s}]=\{\mu\mathbf{o}+(1-\mu)\mathbf{s}:0\leq\mu\leq 1\}$ is contained in $S$. In particular, a set is convex it is star-convex with respect to every $\mathbf{o}\in S$.
[…]
A set $C\subseteq\mathbb{R}^{d}$ is a polyhedral cone if
$C\ =\ \left\{\sum_{i=1}^{k}\mu_{i}\mathbf{v}_{i}:\mu_{1},\ldots,\mu_{k}\in\mathbb{R}_{\geq 0}\right\}$
@@
[…]
A hyperplane $\{\mathbf{x}\in\mathbb{R}^{d}:\langle\mathbf{h},\mathbf{x}\rangle=a\}$ is a supporting hyperplane of the cone $C$ if $\langle\mathbf{h},\mathbf{v}\rangle\geq a$ for all points $\mathbf{v}\in C$. A subset $F\subseteq C$ is called a face if $F=C\cap H$ for some supporting hyperplane $H$.
[…]
@@
Intuitively speaking, a fan is a collection of cones that fit together nicely. Examples of two well-known classes of simplicial fans, namely kite fans and Coxeter fans of type B, are given in Section A.1.
@@
Indeed, for any $\mathbf{x}\in\mathbb{R}^{d}$ there is a unique cone $C\in\Delta$ with generators $\mathbf{v}_{i_{1}}\ldots,\mathbf{v}_{i_{k}}$ such that $\mathbf{x}=\mu_{i_{1}}\mathbf{v}_{i_{1}}+\ldots+\mu_{i_{k}}\mathbf{v}_{i_{k}}\in C$ and $\mu_{i_{j}}>0$ for all $1\leq j\leq k$. If $C$ is a full-dimensional cone, then $k=d$ and $V_{C}=\left(\mathbf{v}_{i_{1}}\dots\mathbf{v}_{i_{d}}\right)$ is an invertible square matrix such that $V_{C}\left(\mu_{i_{1}}\dots\mu_{i_{d}}\right)^{T}=\mathbf{x}$, so $V_{C}^{-1}\mathbf{x}=\left(\mu_{i_{1}}\dots\mu_{i_{d}}\right)^{T}$. Define $\mu_{j}:=0$ for $j\in\{1,\dots,n\}\setminus\{i_{1},\dots,i_{k}\}$. We write $[\mathbf{x}]^{\Delta}\in\mathbb{R}^{n}$ for the vector $(\mu_{1},\dots,\mu_{n})$ expressing $\mathbf{x}$ as a positive linear combination of the generators of the cone of $\Delta$ that it lies in. We will sometimes simply write $[\mathbf{x}]$ when the fan $\Delta$ is clear. For exemplifying computations of $[\mathbf{x}]^{\Delta}$ when $\Delta$ is the kite fan or the Coxeter fan of type B, we refer to Section A.1. Since $f_{\mathbf{a}}^{\Delta}(\mathbf{v}_{i_{j}})=a_{i_{j}}$ for $1\leq j\leq k$, the linearity of $f^{\Delta}_{\mathbf{a}}|_{C}$ implies
@@
Let $X=\{(\mathbf{x}^{(i)},y^{(i)})\}_{i=1}^{m}\subset\mathbb{R}^{d}\times\{0,1\}$ be a binary labeled dataset. Define the $(m\times n)$-matrix $A_{X}$ to be such that the $i$th row is $[\mathbf{x}^{(i)}]^{\Delta}$. Then evaluating $A_{X}\mathbf{a}$ results in a vector whose $i$th entry is $f_{\mathbf{a}}^{\Delta}(\mathbf{x}^{(i)})$. Observe that since $\Delta$ is simplicial, the matrix $A_{X}$ is sparse in the sense that there are at most $d$ non-zero entries in every row.
@@
The classification according to $c_{\mathbf{a}}$ is the vector $(c_{\mathbf{a}}(\mathbf{x}^{(1)}),\dots,c_{\mathbf{a}}(\mathbf{x}^{(m)}))$. By slight abuse of notation we also denote the set of all classifiers $c_{\mathbf{a}}$, $\mathbf{a}\geq\mathbf{0}$ by $S^{\Delta}$. The 00-class $c_{\mathbf{a}}^{-1}(0)$ is enclosed in a star-shaped set. Indeed, it is the union of simplices with vertex sets of the form $\{0,\frac{1}{a_{i_{1}}}\mathbf{v}_{i_{1}},\dots,\frac{1}{a_{i_{k}}}\mathbf{v}_{i_{k}}\}$ where $\mathbf{v}_{i_{1}},\dots,\mathbf{v}_{i_{k}}$ are the generators of a cone $C$ of $\Delta$. We call this a star and denote it as $\operatorname{star}(\mathbf{a})=c_{\mathbf{a}}^{-1}(0)$. See Figure 1 for examples.
@@ 2.2 Loss functions
In this article, we consider minimization with respect to two distinct loss functions: the $0/1$-loss and an exponential loss function. For the $0/1$-loss (or discrete loss), we seek to minimize the number of misclassifications, counting both the false positives and false negatives, i.e.
@@
For the exponential loss function, let $y$ be the random variable giving the class label of the random vector $\mathbf{x}\in\mathbb{R}^{d}$. We approximate the probability that $\mathbf{x}$ is not in the star with the cumulative distribution function of the exponential probability distribution,
$P(y=1|\mathbf{x},\Delta,\mathbf{a})$ = $=\ 1-e^{-\lambda f_{\mathbf{a}}^{\Delta}(\mathbf{x})}\,,$
@@
We observe that $P(y=1|\mathbf{x},\Delta,\mathbf{a})$ approaches 00 when $\mathbf{x}$ approaches 00, i.e. is in the set $\operatorname{star}(\mathbf{a})$ defined by $c_{\mathbf{a}}$, and $1$ when $\mathbf{x}$ approaches infinity, i.e. is outside the star. Furthermore, $\log\left(1-e^{-\lambda f_{\mathbf{a}}^{\Delta}(\mathbf{x}^{(i)})}\right)$ decays exponentially for $f_{\mathbf{a}}^{\Delta}(\mathbf{x}^{(i)})\rightarrow 0$ while $(-\lambda)f_{\mathbf{a}}^{\Delta}(\mathbf{x}^{(i)})$ decreases just linearly for $f_{\mathbf{a}}^{\Delta}(\mathbf{x}^{(i)})\rightarrow\infty$. In particular, (2), which computes the maximum likelihood estimator, has a strong bias to avoid false negatives, while the penalty for false positives is just linear. In contrast, the $0/1$-loss treats both types of misclassifications equally.
@@ 3 Geometry of the parameter space
Let $\Delta$ be a simplicial fan with $n$ generators. Then the VC-dimension of the set of classifiers $S^{\Delta}$ is equal to $n$.
@@
Figure 2: An example of a $1$-dimensional dataset, perfectly classified by a star supported on a fan with $n=2$ rays, and the level sets of the two loss functions in parameter space >$\mathbb{R}^{2}_{>0}$. This example is explained in detail in Section A.2.

### v2

@@ Abstract
We consider binary classification restricted to a class of continuous piecewise linear functions whose decision boundaries are (possibly nonconvex) starshaped polyhedral sets, supported on a fixed polyhedral simplicial fan. We investigate the expressivity of these function classes and describe the combinatorial and geometric structure of the loss landscape, most prominently the sublevel sets, for two loss-functions: the 0/1-loss (discrete loss) and a log-likelihood loss function. In particular, we give explicit bounds on the VC dimension of this model, and concretely describe the sublevel sets of the discrete loss as chambers in a hyperplane arrangement. For the log-likelihood loss, we give sufficient conditions for the optimum to be unique, and describe the geometry of the optimum when varying the rate parameter of the underlying exponential probability distribution.
@@ 1 Introduction
We study the problem of binary classification from a geometric and combinatorial perspective. Given a finite labeled data-set and a prescribed loss-function, we focus on characterizing the structure of those parameters that yield perfect classification – namely, the set of global minimizers of the loss function. More generally, we investigate the geometry and combinatorics of the entire loss landscape in parameter space. Understanding the geometry is central to analyze the behavior of learning algorithms, as, for example, the arrangement of critical points and the connectivity of minimizers influence optimization efficiency and generalization. Combinatorial structures, such as polyhedral decompositions, provide insights into how parameter spaces partition into regions of similar behavior. Naturally, these subdivisions interact with the subdivision into sets of classifiers which induce the same classification on the data, and is therefore intimately related to the VC dimension (the Vapnik-Chervonenkis dimension) of binary classifiers (doi:10.1137/1116025).
@@
In order to make rigorous statements, we fix the function class used for classification to be a class which is suitable for the specific learning task. Fixing a large set of classifiers can lead to practical difficulties due to the complexity of the space of allowed classifiers, while a small set of classifiers may not be able to capture the nature of the underlying problem. A natural function class consists of those functions whose decision boundary – the geometric object separating the two classes – is the boundary of a convex polyhedron. Such function classes have been previously considered, for example, in astorino02_polyhedralseparabilitysuccessive; manwani10_learningpolyhedral and kantchelian2014large, where the optimal separating convex polyhedron is found through iteratively solving LPs, minimizing a logistic loss function and finding a large margin convex separator, respectively.
@@
While polyhedral classifiers form a well-structured function class, they are also highly restrictive. In particular, the region enclosed by the decision boundary is necessarily convex, which may not always align with the structure of the underlying classification problem. Maintaining the piecewise linear nature, but allowing non-convex functions, we consider a class of piecewise linear functions whose decision boundaries are (possibly nonconvex) star-shaped polyhedral sets, supported on a fixed polyhedral fan. Fixing the polyhedral fan makes this a tractable and learnable class of functions whose space of parameters exhibits nice geometric structures as we show. Considering these function classes generalizes the approach in cevikalp2017polyhedral, where kites are used for solving visual object detection and multi-class discrimination.
Our family of star-shaped classifiers falls within the broader class of continuous piecewise-linear functions, and as such can be represented by a suitably structured ReLU network. However, the powerful flexibility of ReLU networks makes it challenging to enforce specific geometric properties, such as ensuring that the decision region satisfies star-convexity. Moreover, the parameter space of general ReLU neural networks admits undesirable combinatorial and geometric properties such as disconnectedness and local non-global minima even for separable data brandenburg2024the. In contrast, by building our classifiers directly on a fixed simplicial fan, we retain the ability to model non-convex boundaries and yet maintain high control over the shape and connectivity of the decision regions.
@@
We focus on classification with polyhedral starshaped sets with respect to two loss functions: the 0/1-loss (or discrete loss) allows us to study the parameter space in a combinatorial fashion using polyhedral methods. Additionally, we consider a log-likelihood loss function, which is amenable to numerical optimization methods, and whose level sets carry convex-geometric structures.
@@ 1.1 Our contributions
We initiate the geometric study of the space of parameters defining the classification by polyhedral starshaped sets, and show that binary classifications with this model correspond to chambers in the data arrangement, a hyperplane arrangement within the parameter space. Investigating the expressivity of such starshaped polyhedral classifiers, we show that the VC dimension equals the number of rays in the fan, quantifying how the number of linear regions of the classifier impacts sample complexity.
@@
We examine the effect of the choice of the loss function explicitly on two concrete loss functions, and contrast how the geometry induced by these different losses governs optimization. We show that the sublevel sets of false positives and false negatives are starshaped sets in the parameter space, and, in case of perfect separability, sublevel sets with respect to the discrete loss are starconvex sets. For the log-likelihood loss we show concavity, and consequently convexity holds for its superlevel sets, implying that an optimum can be found in polynomial time. We give sufficient conditions for the optimum to be unique, and describe the geometry of the optimum when varying the rate parameter of the underlying exponential probability distribution.
@@
While most of our results focus on starshaped sets with fixed origin, we also consider starshaped classifiers where we allow translation of the origin, and consider the translation vector as an extra parameter. For a fixed starshaped set and varying translation vectors, we show that the discrete loss is constant on chambers in an arrangement of stars, but the sublevel sets are generally no longer starconvex. While this setup combined with the log-likelihood loss does in general not lead to a convex program, we show that the log-likelihood function is piecewise concave on the underlying data fan arrangement.
@@ 1.2 Limitations
Throughout the article, we assume a fixed simplicial fan as given. In practical applications, an appropriate fan suitable for the specific task needs to be chosen prior to the analysis. We emphasize that this paper is a purely theoretical contribution. The presented framework has not been tested on large-scale synthetic or real-world data, only small-scale experiments as presented in the end of this article have been conducted. Developing a systematic or heuristic approach for selecting parameters such as the simplicial fan, a translation vector or the rate parameter, is beyond the scope of this work.
@@ 2.1 Polyhedral geometry and stars
We begin by introducing essential notions from polyhedral geometry, and the class of classifiers we consider. For a thorough background on polyhedral geometry we refer the reader to (ziegler_lecturespolytopes, Chapters 1-2). Examples and visualizations of polyhedral fans are given in Appendix˜A
Definition 2.1.
A set $S\subseteq\mathbb{R}^{d}$ is star-convex with respect to a center $\mathbf{o}\in S$ if for every $\mathbf{s}\in S$, the line segment $\mathbf{o},\mathbf{s}]=\{\mu\mathbf{o}+(1-\mu)\mathbf{s}:0\leq\mu\leq 1\}$ is contained in $S$. In particular, a set is convex it is star-convex with respect to every $\mathbf{o}\in S$.
Definition 2.2.
A set $C\subseteq\mathbb{R}^{d}$ is a polyhedral cone if
@@
Definition 2.3.
A hyperplane $\{\mathbf{x}\in\mathbb{R}^{d}:\langle\mathbf{h},\mathbf{x}\rangle=a\}$ is a supporting hyperplane of the cone $C$ if $\langle\mathbf{h},\mathbf{v}\rangle\geq a$ for all points $\mathbf{v}\in C$. A subset $F\subseteq C$ is called a (proper) face if $F=C\cap H$ for some supporting hyperplane $H$.
Definition 2.4.
@@
Intuitively speaking, a fan is a collection of cones that fit together nicely. Examples of two well-known classes of simplicial fans, namely kite fans and Coxeter fans of type B, are given in Section˜A.1.
@@
Indeed, for any $\mathbf{x}\in\mathbb{R}^{d}$ there is a unique cone $C\in\Delta$ with generators $\mathbf{v}_{i_{1}}\ldots,\mathbf{v}_{i_{k}}$ such that $\mathbf{x}=\mu_{i_{1}}\mathbf{v}_{i_{1}}+\ldots+\mu_{i_{k}}\mathbf{v}_{i_{k}}\in C$ and $\mu_{i_{j}}>0$ for all $1\leq j\leq k$. If $C$ is a full-dimensional cone, then $k=d$ and $V_{C}=\left(\mathbf{v}_{i_{1}}\dots\mathbf{v}_{i_{d}}\right)$ is an invertible square matrix such that $V_{C}\left(\mu_{i_{1}}\dots\mu_{i_{d}}\right)^{T}=\mathbf{x}$, so $V_{C}^{-1}\mathbf{x}=\left(\mu_{i_{1}}\dots\mu_{i_{d}}\right)^{T}$. Define $\mu_{j}:=0$ for $j\in\{1,\dots,n\}\setminus\{i_{1},\dots,i_{k}\}$. We write $\mathbf{x}]^{\Delta}\in\mathbb{R}^{n}$ for the vector $(\mu_{1},\dots,\mu_{n})$ expressing $\mathbf{x}$ as a positive linear combination of the generators of the cone of $\Delta$ that it lies in. We will sometimes simply write $\mathbf{x}]$ when the fan $\Delta$ is clear. For exemplifying computations of $\mathbf{x}]^{\Delta}$ when $\Delta$ is the kite fan or the Coxeter fan of type B, we refer to Section˜A.1. Since $f_{\mathbf{a}}^{\Delta}(\mathbf{v}_{i_{j}})=a_{i_{j}}$ for $1\leq j\leq k$, the linearity of $f^{\Delta}_{\mathbf{a}}|_{C}$ implies
@@
Let $X=\{(\mathbf{x}^{(i)},y^{(i)})\}_{i=1}^{m}\subset\mathbb{R}^{d}\times\{0,1\}$ be a binary labeled dataset. Define the $(m\times n)$-matrix $A_{X}$ to be such that the $i$th row is $\mathbf{x}^{(i)}]^{\Delta}$. Then evaluating $A_{X}\mathbf{a}$ results in a vector whose $i$th entry is $f_{\mathbf{a}}^{\Delta}(\mathbf{x}^{(i)})$. Observe that since $\Delta$ is simplicial, the matrix $A_{X}$ is sparse in the sense that there are at most $d$ non-zero entries in every row.
@@
The classification according to $c_{\mathbf{a}}$ is the vector $(c_{\mathbf{a}}(\mathbf{x}^{(1)}),\dots,c_{\mathbf{a}}(\mathbf{x}^{(m)}))$. By slight abuse of notation we also denote the set of all classifiers $c_{\mathbf{a}}$, $\mathbf{a}\geq\mathbf{0}$ by $S^{\Delta}$. The 0-class $c_{\mathbf{a}}^{-1}(0)$ is enclosed in a star-shaped set. Indeed, it is the union of simplices with vertex sets of the form $\{0,\frac{1}{a_{i_{1}}}\mathbf{v}_{i_{1}},\dots,\frac{1}{a_{i_{k}}}\mathbf{v}_{i_{k}}\}$ where $\mathbf{v}_{i_{1}},\dots,\mathbf{v}_{i_{k}}$ are the generators of a cone $C$ of $\Delta$. We call this a star and denote it as $\operatorname{star}(\mathbf{a})=c_{\mathbf{a}}^{-1}(0)$. See Figure 1 for examples.
@@ 2.2 Loss functions
In this article, we consider minimization with respect to two distinct loss functions: the $0/1$-loss and a log-likelihood loss function. For the $0/1$-loss (or discrete loss), we seek to minimize the number of misclassifications, counting both the false positives and false negatives, i.e.
@@
For the log-likelihood loss function, let $y$ be the random variable giving the class label of the random vector $\mathbf{x}\in\mathbb{R}^{d}$. We approximate the probability that $\mathbf{x}$ is not in the star with the cumulative distribution function of the exponential probability distribution,
$P(y=1|\mathbf{x},\Delta,\mathbf{a})\$= $=\ 1-e^{-\lambda f_{\mathbf{a}}^{\Delta}(\mathbf{x})}\,$,
@@
We observe that $P(y=1|\mathbf{x},\Delta,\mathbf{a})$ approaches 0 when $\mathbf{x}$ approaches 0, i.e. is in the set $\operatorname{star}(\mathbf{a})$ defined by $c_{\mathbf{a}}$, and $1$ when $\mathbf{x}$ approaches infinity, i.e. is outside the star.
@@ 3 Geometry of the parameter space
Let $\Delta$ be a simplicial fan with $n$ generators. Then the VC dimension of the set of classifiers $S^{\Delta}$ is equal to $n$.
@@
Figure 2: An example of a $1$-dimensional dataset, perfectly classified by a star supported on a fan with $n=2$ rays, and the level sets of the two loss functions in parameter space $\mathbb{R}^{2}_{>0}$. This example is explained in detail in Section˜A.2.